# Asymmetrical thermonuclear supernovae triggered by the tidal disruption of white dwarfs

### v1

@@ 4.2 Results for the 1D-equivalent models computed with CMFGEN
Figure 9: Radiative properties of the 1D counterpart calculated with CMFGEN and based on the 3D models computed with AREPO. We show the bolometric light curves (left) and the maximum-light optical spectra for our model set (right).
@@
The spectral evolution computed with CMFGEN for the $b=0.15$, $0.17$, $0.18$, and $0.19$ models at $1.5$, 55, $10$, 2020 and $45\,\mathrm{d}$ is shown in Figure 10. The $b=0.10$ and $b=0.12$ models with larger $^{56}\rm{N}$i masses than the $b=0.15$ model have a similar spectral morphology and a global offset in flux (thus not shown). Most of the diversity occurs at the lower end of the $^{56}\rm{N}$i distribution where the contribution from IMEs strengthens. This is made more evident in Figure 11 where all models computed with CMFGEN are shown at $100\,\mathrm{d}$. The difference at this later time is the additional presence of forbidden lines, primarily of [Co iii] $5888.5\,\mathrm{\AA }$ and [Fe iii] $4658.0\,\mathrm{\AA }$ in models with a high $^{56}\rm{N}$i mass whereas [O i] $6316.0\,\mathrm{\AA }$ and [Ca ii] $7307.7\,\mathrm{\AA }$ dominate in our models with modest amounts of $^{56}\rm{N}$i. Hence, this grid effectively covers from SNe Ia ejecta entirely dominated by IGEs like 91T-like events (for a review, see Phillips et al. 2024), to standard Type Ia SN ejecta (e.g., SN 2002bo; Benetti et al. 2004; Blondin et al. 2015), and down to ejecta that are more similar in composition to Type Ic SNe (C and O dominated with little $^{56}\rm{N}$i; see, e.g., Nomoto et al. 1994).
This evolution in time and relative to $^{56}\rm{N}$i mass is similar to what has been shown in numerous other studies for the explosion of both Chandrasekhar and sub-Chandrasekhar white dwarfs (see, e.g., Hoeflich and Khokhlov 1996; Sim et al. 2010; Blondin et al. 2013; Pakmor et al. 2013; Dessart et al. 2014) and does not warrant further discussion. It is done here to set the stage for the later discussion on the 2D simulations.
Figure 11: Montage of spectra for the D-equivalent of the 3D models computed with CMFGEN at $100\,\mathrm{d}$ after the tidal disruption. A color coding is used to differentiate the emission from O, Ca, Co, and Fe, which dominate at that late time.
@@ 5 Results from 2D radiative-transfer calculations
The 1D simulations presented in the preceding section were post-processed at multiple epochs with the 2D, steady-state polarized radiative-transfer code LONG_POL. In practice, the 2D code reads the 1D wavelength-dependent total (line and continuum) opacities $\chi$, emissivities $\eta$, and electron densities $N_{e}$ versus the velocities output by the CMFGEN calculations and remaps these onto the 2D, spherical polar grid of LONG_POL. The asymmetry of the 3D AREPO simulations was encoded in LONG_POL through a latitudinal scaling of $\chi$, $\eta$, and $N_{e}$. Below a polar angle of $\theta_{0}$ (relative to the axis of the cone at $\theta=0$ deg), this scaling factor was set to unity, whereas beyond it was set to vary as $\alpha+(1-\alpha)\exp(-\beta^{2})$, where $\beta=(\theta-\theta_{0})/\Delta\theta$. To reflect the asymmetry of the 3D debris, we adopted $\alpha=10^{-8}$, $\theta_{0}=15\deg$, and $\Delta\theta=10\deg$. We found that the value of $\alpha$ had to be very low in order to quench the contribution from the material surrounding the unbound ejecta, particularly at early times when the debris is still dense and optically thick.
Obviously, this treatment of the asymmetry is very rough. When assuming spherical symmetry, the deeper ejecta layers are not only denser, which leads to a smaller photon mean free path, but they are also enshrouded within the overlying ejecta layers, which may contribute a significant optical depth. This affects the propagation of both $\gamma$-rays (i.e., from radioactive decay emission) and low-energy photons. When one allows for asymmetry (whether truly in 3D or in the approximate morphology in 2D), the deeper layers are equally dense as in the 1D case, but they are no longer enshrouded. This holds for the sides of the ‘cone’, or for the innermost layers which are effectively bare. In our post-treatment of the 1D CMFGEN simulations with LONG_POL, we ignored this aspect and thus overestimate the temperature and ionization of these deeper layers. This means that we overestimate the emissivity of the deeper layers, which affects all observables for viewing angles away from $\theta=0\deg$. More specifically, we overestimate the brightness (a global shift in all bands), the color (predictions will be overly blue in the optical), and the ionization (we may predict emission from twice-ionized species, whereas once-ionized species may dominate). This issue is, however, not as bad as it may seem since the inner ejecta are $10^{5}$ times denser than the outer, fast-moving ejecta regions. Thus, even if the inner ejecta regions are bare, the mean-free-path of both low- and high-energy photons is small, and radiation is considerably more trapped in the inner ejecta regions than in the outer ejecta regions. A more suitable approach would be to solve the full radiative transfer of low- and high-energy photons in 3D, as done in (MacLeod et al., 2016) – this is deferred to future work.
@@
Figure 12: Photometric properties of SNe Ia TDE models computed with LONG_POL. Top left: $B$-band light curve of the $b=0.16$ model for viewing angles of 0, $90$, and $180\deg$. We also overplot the CMFGEN results (assuming spherical symmetry). Other panels: Same as top left, but comparing different models for a given viewing angle.
@@
Figure 13 shows the spectral evolution from about 5 to $40\,\mathrm{d}$ for the $b=0.16$ model for a viewing angle of 0, $90$, and $180\deg$. This covers most of the photospheric phase and thus straddles the time of bolometric maximum. Hence, the spectra shift upwards until about $20\,\mathrm{d}$ before shifting downwards towards a lower power. For a viewing angle of $0\deg$, one observes the debris along the direction of expansion. At the earliest time, the spectra show much structure and rapidly change with wavelength. This arises from the fact that we are essentially observing the troughs of what would in 1D appear as a fully developed P-Cygni profile with both absorption and a broad emission (see, e.g., Figure 9 or Figure 10) – only the absorption part is essentially present for that viewing angle because the debris are confined within a narrow cone pointing towards the observer. As time progresses, the photosphere recedes in the ejecta, lines become narrower but still exhibit a systematic blueshift.
@@
Figure 15 is a counterpart of Figure 14 but now at $100\,\mathrm{d}$. At nebular conditions, the spectrum is dominated by a few strong emission lines, including gaps with no emission beyond $5500\,\mathrm{\AA }$, essentially associated with regions devoid of Fe ii lines. Here too, some insets are used to reveal the wavelength shifts in some strong emission lines, using as diagnostics the air wavelength of [Fe iii] $4658\,\mathrm{\AA }$, the gfgf-weighted333gfgf refers to the oscillator strength of an atomic transition. mean wavelength of the Ca ii NIR triplet, or that of [Ca ii] $\lambda\lambda$7291, 73237291,\,7323.
@@
Figure 13: Spectral evolution in the optical for the 2D $b=0.16$ model and for viewing angles of $0\deg$ (top), $90\deg$ (middle), and $180\deg$ (bottom).
Figure 14: Optical spectra for models $b=0.10$ (top), $b=0.16$ (middle), and $b=0.19$ (bottom) at the time of bolometric maximum (of the 1D model counterpart) for viewing angles of 0, $90$, and $180\deg$.
Figure 15: Optical spectra for models $b=0.10$ (top), $b=0.16$ (middle), and $b=0.19$ (bottom) at $100\,\mathrm{d}$ and for viewing angles of 0, $90$, and $180\deg$.
@@ 6 Discussion and conclusions
SNe Ia TDEs are most dramatically distinct from garden-variety SNe Ia in the extreme asymmetry of the debris and the presence of a central BH. Even if assumed dormant, the gravitational pull on the ejecta extends the dynamical phase (i.e., the debris moves ballistically only after a number of hours) and leads to a fallback that carves the inner ejecta below a few $1000\,\mathrm{km}\,\mathrm{s}^{-1}$. The debris asymmetry also leads to a dependence of brightness and colors with viewing angles, although this also depends on the $^{56}\rm{N}$i mass and the chemical stratification. Although not discussed in this work, our 2D radiative-transfer calculations predict a polarization at the 1% level around bolometric maximum, thus well above any polarization measured for the more spherical, standard Type Ia SNe (see, for example, Wang and Wheeler 2008). The most striking spectral signature of SNe Ia TDEs, best visible at nebular times, is the large velocity shifts of emission lines for viewing angles away from $90\deg$, which arise from the strong ejecta asymmetry. This shift is aggravated by the cavity in the inner ejecta, resulting in a large displacement of the line emission (i.e., both a skewness and a shift of line emission). Interestingly, this has been observed in a fast-blue optical transient with systematically blueshifted and displaced H i emission lines (Gutiérrez et al., 2024) – this event was interpreted as the TDE of an unevolved, low-mass star by an IMBH.
@@ Acknowledgements.
PV thanks Bart Ripperda and Norman Murray for providing access to their computational time on Compute Canada’s Niagara cluster. LD acknowledges support from the ESO Scientific Visitor Program for a visit to ESO-Garching during the summer of 2025.
Facilities. Compute Canada’s Niagara cluster, CITA’s Sunnyvale cluster. This work was granted access to the HPC resources of TGCC under the allocation 2024 – A0170410554 on Irene-Rome made by GENCI, France.
Software. AREPO (Springel, 2010; Pakmor et al., 2016; Weinberger et al., 2020), V1D (Livne, 1993; Dessart et al., 2010), CMFGEN (Hillier and Miller, 1998; Hillier and Dessart, 2012), LONG_POL (Hillier, 1994, 1996; Dessart and Hillier, 2011; Dessart et al., 2021)
Author Contributions. PV and LD contributed equally to the paper, with them conducting, analyzing, and writing the hydrodynamics simulations and the radiative transfer calculations, respectively. All authors contributed to the project planning, analysis, and writing.

### v2

@@ 4.2 Results for the 1D-equivalent models computed with CMFGEN
Figure 9: Radiative properties of the 1D counterpart calculated with CMFGEN and based on the 3D models computed with AREPO. We show the bolometric light curves (left) and the maximum-light optical spectra for our model set (right). Closer encounters (e.g., $b=0.10$) are brighter and show stronger Fe iii lines in their spectra, while wider encounters (e.g., $b=0.19$) are dimmer and show stronger O i, Si $i$, and Ca ii lines in their spectra.
Figure 11: Montage of spectra for the 1D-equivalent of the 3D models computed with CMFGEN at $100\,\mathrm{d}$ after the tidal disruption. A color coding is used to differentiate the emission from O, Ca, Co, and Fe, which dominate at that late time. The spectra of closer encounters with larger $^{56}\rm{N}$i masses (e.g., $b=0.10$) are dominated by Co iii and Fe iii lines, while those of wider encounters with lower $^{56}\rm{N}$i masses (e.g., $b=0.19)$ are dominated by O i and Ca ii lines.
@@
The spectral evolution computed with CMFGEN for the $b=0.15$, $0.17$, $0.18$, and $0.19$ models at $1.5$, 55, $10$, 2020 and $45\,\mathrm{d}$ is shown in Figure 10. The $b=0.10$ and $b=0.12$ models with larger $^{56}\rm{N}$i masses than the $b=0.15$ model have a similar spectral morphology and a global offset in flux (thus not shown). Most of the diversity occurs at the lower end of the $^{56}\rm{N}$i distribution where the contribution from IMEs strengthens. This is made more evident in Figure 11 where all models computed with CMFGEN are shown at $100\,\mathrm{d}$. The difference at this later time is the additional presence of forbidden lines, primarily of [Co iii] $5888.5\,\mathrm{\AA }$ and [Fe iii] $4658.0\,\mathrm{\AA }$ in models with a high $^{56}\rm{N}$i mass whereas [O i] $6316.0\,\mathrm{\AA }$ and [Ca ii] $7307.7\,\mathrm{\AA }$ dominate in our models with modest amounts of $^{56}\rm{N}$i. Hence, this grid effectively covers from SNe Ia ejecta entirely dominated by IGEs similar to 91T-like events (for a review, see Phillips et al. 2024), to standard Type Ia SN ejecta (e.g., SN 2002bo; Benetti et al. 2004; Blondin et al. 2015), and down to ejecta that are more similar in composition to Type Ic SNe (C and O dominated with little $^{56}\rm{N}$i; e.g., Nomoto et al. 1994).
This evolution in time and relative to $^{56}\rm{N}$i mass is similar to what has been shown in numerous other studies for the explosion of both Chandrasekhar and sub-Chandrasekhar white dwarfs (e.g., Hoeflich and Khokhlov 1996; Sim et al. 2010; Blondin et al. 2013; Pakmor et al. 2013; Dessart et al. 2014) and does not warrant further discussion. It is done here to set the stage for the later discussion on the 2D simulations.
@@ 5 Results from 2D radiative-transfer calculations
Figure 12: Photometric properties of SNe Ia TDE models computed with LONG_POL. Top left: $B$-band light curve of the $b=0.16$ model for viewing angles of 0, $90$, and $180\deg$. We also overplot the CMFGEN results (assuming spherical symmetry). Other panels: Same as top left, but comparing different models for a given viewing angle. The $B$-band brightness is greater for closer encounters and viewing angles away from $0\deg$.
The 1D simulations presented in the preceding section were post-processed at multiple epochs with the 2D, steady-state polarized radiative-transfer code LONG_POL. In practice, the 2D code reads the 1D wavelength-dependent total (line and continuum) opacities $\chi$, emissivities $\eta$, and electron densities $N_{e}$ versus the velocities output by the CMFGEN calculations and remaps these onto the 2D, spherical polar grid of LONG_POL. The asymmetry of the 3D AREPO simulations was encoded in LONG_POL through a latitudinal scaling of $\chi$, $\eta$, and $N_{e}$. Below a polar angle of $\theta_{0}$ (relative to the axis of the cone at $\theta=0$ deg), this scaling factor was set to unity, whereas beyond it was set to vary as $\alpha+(1-\alpha)\exp(-\beta^{2})$, where $\beta=(\theta-\theta_{0})/\Delta\theta$. To reflect the asymmetry of the 3D debris, we adopted $\alpha=10^{-8}$, $\theta_{0}=15\deg$, and $\Delta\theta=10\deg$. We found that the value of $\alpha$ had to be very low in order to quench the contribution from the material surrounding the unbound ejecta, particularly at early times when the debris is still dense and optically thick. Figure 18 illustrates this latitudinal scaling.
Obviously, this treatment of the asymmetry is very rough. When assuming spherical symmetry, the deeper ejecta layers are not only denser, which leads to a smaller photon mean free path, but they are also enshrouded within the overlying ejecta layers, which may contribute a significant optical depth. This affects the propagation of both $\gamma$-rays (i.e., from radioactive decay emission) and low-energy photons. When one allows for asymmetry (whether truly in 3D or in the approximate morphology in 2D), the deeper layers are equally dense as in the 1D case, but they are no longer enshrouded. This holds for the sides of the “cone”, or for the innermost layers which are effectively bare. In our post-treatment of the 1D CMFGEN simulations with LONG_POL, we ignored this aspect and thus overestimate the temperature and ionization of these deeper layers. This means that we overestimate the emissivity of the deeper layers, which affects all observables for viewing angles away from $\theta=0\deg$. More specifically, we overestimate the brightness (a global shift in all bands), the color (predictions will be overly blue in the optical), and the ionization (we may predict emission from twice-ionized species, whereas once-ionized species may dominate). This issue is, however, not as bad as it may seem since the inner ejecta are $10^{5}$ times denser than the outer, fast-moving ejecta regions. Thus, even if the inner ejecta regions are bare, the mean-free-path of both low- and high-energy photons is small, and radiation is considerably more trapped in the inner ejecta regions than in the outer ejecta regions. A more suitable approach would be to solve the full radiative transfer of low- and high-energy photons in 3D, as done in MacLeod et al. (2016) – this is deferred to future work.
@@
Figure 13: Spectral evolution in the optical for the 2D $b=0.16$ model and for viewing angles of $0\deg$ (top), $90\deg$ (middle), and $180\deg$ (bottom). The spectra change rapidly with wavelength at earlier times and lines become narrower with time. The spectral lines at $0\deg$ and $180\deg$ are blueshifted and redshifted, respectively (most evident at later times).
Figure 14: Optical spectra for models $b=0.10$ (top), $b=0.16$ (middle), and $b=0.19$ (bottom) at the time of bolometric maximum (of the 1D model counterpart) for viewing angles of 0, $90$, and $180\deg$. The zoomed-in panels show wavelength shifts in the spectral lines that depend on viewing angle.
Figure 15: Optical spectra for models $b=0.10$ (top), $b=0.16$ (middle), and $b=0.19$ (bottom) at $100\,\mathrm{d}$ and for viewing angles of 0, $90$, and $180\deg$. The zoomed-in panels show wavelength shifts in the spectral lines that depend on viewing angle.
@@
Figure 13 shows the spectral evolution from about 5 to $40\,\mathrm{d}$ for the $b=0.16$ model for a viewing angle of 0, $90$, and $180\deg$. This covers most of the photospheric phase and thus straddles the time of bolometric maximum. Hence, the spectra shift upwards until about $20\,\mathrm{d}$ before shifting downwards towards a lower power. For a viewing angle of $0\deg$, one observes the debris along the direction of expansion. At the earliest time, the spectra show much structure and rapidly change with wavelength. This arises from the fact that we are essentially observing the troughs of what would in 1D appear as a fully developed P-Cygni profile with both absorption and a broad emission (see Figure 9 or Figure 10) – only the absorption part is essentially present for that viewing angle because the debris are confined within a narrow cone pointing towards the observer. As time progresses, the photosphere recedes in the ejecta, lines become narrower but still exhibit a systematic blueshift.
@@
Figure 15 is a counterpart of Figure 14 but now at $100\,\mathrm{d}$. At nebular conditions, the spectrum is dominated by a few strong emission lines, including gaps with no emission beyond $5500\,\mathrm{\AA }$, essentially associated with regions devoid of Fe ii lines. Here too, some insets are used to reveal the wavelength shifts in some strong emission lines, using as diagnostics the air wavelength of [Fe iii] $4658\,\mathrm{\AA }$, the gfgf-weighted555gfgf refers to the oscillator strength of an atomic transition. mean wavelength of the Ca ii NIR triplet, or that of [Ca ii] $\lambda\lambda$7291, 73237291,\,7323.
@@ 6 Discussion and conclusions
SNe Ia TDEs are most dramatically distinct from garden-variety SNe Ia in the extreme asymmetry of the debris and the presence of a central BH. Even if assumed dormant, the gravitational pull on the ejecta extends the dynamical phase (i.e., the debris moves ballistically only after a number of hours) and leads to a fallback that arves the inner ejecta below a few $1000\,\mathrm{km}\,\mathrm{s}^{-1}$. The debris asymmetry also leads to a dependence of brightness and colors with viewing angles, although this also depends on the $^{56}\rm{N}$i mass and the chemical stratification. The most striking spectral signature of SNe Ia TDEs, best visible at nebular times, is the large velocity shifts of emission lines for viewing angles away from $90\deg$, which arise from the strong ejecta asymmetry. This shift is aggravated by the cavity in the inner ejecta, resulting in a large displacement of the line emission (i.e., both a skewness and a shift of line emission). Interestingly, this has been observed in a fast-blue optical transient with systematically blueshifted and displaced H i emission lines (Gutiérrez et al., 2024) – this event was interpreted as the TDE of an unevolved, low-mass star by an IMBH.
@@ Acknowledgements.
PV thanks Bart Ripperda and Norman Murray for providing access to their computational time on Compute Canada’s Niagara cluster. PV thanks CITA for access to the Sunnyvale cluster. LD acknowledges support from the ESO Scientific Visitor Program for a visit to ESO-Garching during the summer of 2025. LD acknowledges access to the HPC resources of TGCC under the allocation 2024 – A0170410554 on Irene-Rome made by GENCI, France.
PV and LD contributed equally to the paper, with them conducting, analyzing, and writing the hydrodynamics simulations and the radiative transfer calculations, respectively. All authors contributed to the project planning, analysis, and writing.
Appendix A Density profile of the white dwarf
Figure 16: Density profiles of the AREPO-initialized $0.6\,\mathrm{M}_{\odot}$ WD ($10^{-7}\penalty 10000\ \,\mathrm{M}_{\odot}/\mathrm{cell}$ resolution) as a function of mass (left) and radial (right) coordinates. The blue dashed and the orange solid lines denote the profiles initially (identical to 1D MESA profile) and after 10 sound-crossing times ($\sim 127\,\mathrm{s}$). The profiles overlap significantly.
Appendix B Kinetic energies of the unbound ejecta
Figure 17: Kinetic energies $E_{\mathrm{K,unbound}}$ of the unbound ejecta per 0.5 dex in density at $\sim 200\,\mathrm{s}$ and $\sim 500\,\mathrm{s}$, for $b$ ranging from $0.17$ to $0.20$. For $b=0.20$ (and $b=0.19$), $E_{\mathrm{K,unbound}}$ at $\sim 500\,\mathrm{s}$ is dominated by low-density, (very) high-velocity material, and is hence larger than at $\sim 200\,\mathrm{s}$. However, as $b$ decreases (e.g., $b=0.17$ and lower, not shown) and nuclear fusion becomes significant, there is only a slight change in $E_{\mathrm{K,unbound}}$ with time owing to the decreased presence of such material. The peaks of the kinetic energies shift to lower densities with time due to the expansion of the ejecta.
Appendix C Latitudinal scaling in 2D radiative transfer
Figure 18: Latitudinal scaling used for the inputs (opacities $\chi$, emissivities $\eta$, and electron densities $N_{e}$) to the 2D radiative transfer code LONG_POL. The dashed lines represent the boundaries of a cone of $30^{\circ}$. The parameters $\alpha$ and $\beta$ set the steepness of scaling drop-off beyond the conical region and the gaussian scaling centered about the x-x direction, respectively. This scaling is performed to mimic the asymmetry of the WD TDE ejecta, shown in Figure 1.